# Quantum dust cores of black holes and their quasi-normal modes

### v1

@@ 1 Introduction
Black holes are some of the most challenging physical prediction of General Relativity (GR). Traditionally, they are described by stationary vacuum solutions of the Einstein equations, which are characterised by the presence of an event horizon and by their geodesic incompleteness. Spacetime singularities mark the breakdown of the classical theory of gravity, and there is a growing consensus that they should not show up in a complete, but still unknown, theory of quantum gravity [2]. Moreover, their trivial vacuum structure is altered if one takes into account the matter that collapses in the process of black hole formation.
In this context, Ref. [1] studies a quantum model of black holes that makes the singularity integrable. In this model, a black hole forms from the gravitational collapse of an isotropic distribution of dust, partitioned into nested layers of dust particles moving in the Schwarzschild spacetime
@@
where $m=m(r)$ is the Misner-Sharp-Hernandez (MSH) fraction of the Arnowitt-Deser-Misner (ADM) mass inside a sphere of radius $r=r(\tau)$ and $\mbox{${\rm d}$}\Omega^{2}=\mbox{${\rm d}$}\theta^{2}+\sin^{2}{\theta}\,\mbox{${\rm d}$}\varphi^{2}$ [^1] We shall always use units with $c=1$ and often write the Planck constant $\hbar=\ell_{\rm p}\,m_{\rm p}$ and the Newton constant $G_{\rm N}=\ell_{\rm p}/m_{\rm p}$, where $\ell_{\rm p}$ and $m_{\rm p}$ are the Planck length and mass, respectively. Each dust particle carries the same proper mass $\mu$, while the $i^{\mathrm{th}}$ layer has a MSH mass $\mu_{i}$ for $i=0,\ldots,N$, where $i=0$ refers to the innermost core, and $i=N$ labels the outermost shell. The cumulative mass $M_{i}$ inside the $i^{\rm th}$ layer is then given by
@@
The canonical quantisation prescription leads to a time-independent Schrödinger equation, whose solutions are given by the Hamiltonian eigenfunctions $\psi_{n_{i}}$ corresponding to the eigenvalues
@@
Since $E_{i}^{2}\geq 0$, Eq. (1.3) sets the bound $n_{i}\geq N_{i}$, where  [^2] See also Ref. [3].
@@
[…]
@@
This mass function weakens the central singularity of the Schwarzschild solution and matches the total ADM mass at the surface of the outermost layer, as shown in the left panel of Fig. 1. Eq. (1.11) thus provides an effective description of the geometry of a collapsed core of radius $R_{\rm s}\simeq\bar{R}_{N}+\Delta{R}_{N}\simeq 3\,G_{\rm N}\,M/2$, shorter than, but still comparable to, the gravitational radius $R_{\rm H}=2\,G_{\rm N}\,M$.
@@ 2.1 Quantum mass refinement
The wave-function (1.6) of each layer shows an oscillatory damped pattern that suggests an overlap among (at least) the nearest neighbours. This effect turns the linear mass profile into a (slightly) parabolic mass distribution inside the ball [7]. In fact, particles from the $i^{\rm{th}}$ layer have non-vanishing probability to be located inside the $j^{\rm{th}}$ layer given by Eq. (2.2) (see right panel of Fig. 1).
@@
with $x=r/R_{H}$ and $a$, $b$, $c$ fitting coefficients [7] [^3] The profile does not significantly depend on the number $N$ of layers since the largest faction of the total mass is located in the outer layers. This refined distribution preserves the integrability of the metric in the origin and does not give rise to any inner horizons. Values of $\mathcal{M}_{i}$ and the parabolic profile are shown in the left panel of Fig. 2.
@@ 3 Quasi-normal modes
[…]
@@
and the next goal is to determine the matrix elements $M_{ij}$. The final step consists into matching WKB solutions in region I and III with the one in region II. This is done by Taylor expanding $Q(r_{*})$ around its maximum and solve equation (3.9) with a combination of parabolic cylinder functions. The coefficients of these asymptotically expanded solutions, can be matched with those in the outer region and used to populate the matrix (3.15). The boundary conditions for $\xi$ at $r_{*}\to\pm\infty$ follow from requiring that no waves can escape from the horizon, i.e. $A^{(\mathrm{in})}_{\rm H}=0$, and that there are no incoming waves, i.e. $A^{(\mathrm{in})}_{\infty}=0$ [^5] These arguments are reversed if $\omega_{R}<0$. These conditions on the amplitudes forces the respective coefficients $M_{ij}$ to vanish, which finally allows us to derive the expression
@@
We also remark that the calculations for the interpolating case are performed to fifth order for practical considerations, since higher orders of Padé approximants become more computationally demanding than advantageous. For some values of $n$ and $\ell$ [^6] Specifically when $n\geq\ell$ where the method is less accurate. the polynomials involve very small numbers that compromise the numerical stability of their asymptotic convergence, rather than improving the precision. This is not a real drawback, as fifth-order frequencies are very accurate and there is no a priori reason to force higher-order corrections [14].
@@ 4 Conclusion and outlooks
[…]

### v2

@@ 1 Introduction
Black holes are some of the most challenging physical prediction of General Relativity (GR). Traditionally, they are described by stationary vacuum solutions of the Einstein equations, which are characterised by the presence of an event horizon and by their geodesic incompleteness. Spacetime singularities mark the breakdown of the classical theory of gravity, and there is a growing consensus that they should not show up in a complete, but still unknown, theory of quantum gravity [2]. Moreover, their trivial vacuum structure is likely altered if one takes into account the quantum nature of matter that collapses and forms astrophysical black holes.
Fundamental constituents of matter in the standard model of particle physics are represented by quantum excitations of fields that behave like localised objects (particles) in suitable regimes (when particle number is conserved), and any realistic modelling of black holes should somehow take into account these features. In this context, a quantum model was studied in Ref. [1] in which a black hole forms from the gravitational collapse of an isotropic distribution of dust, partitioned into nested layers of dust particles moving in the Schwarzschild spacetime
@@
where $m=m(r)$ is the Misner-Sharp-Hernandez (MSH) fraction of the Arnowitt-Deser-Misner (ADM) mass inside a sphere of radius $r=r(\tau)$ and $\mbox{${\rm d}$}\Omega^{2}=\mbox{${\rm d}$}\theta^{2}+\sin^{2}{\theta}\,\mbox{${\rm d}$}\varphi^{2}$.  [^1] We shall always use units with $c=1$ and often write the Planck constant $\hbar=\ell_{\rm p}\,m_{\rm p}$ and the Newton constant $G_{\rm N}=\ell_{\rm p}/m_{\rm p}$, where $\ell_{\rm p}$ and $m_{\rm p}$ are the Planck length and mass, respectively. Each dust particle carries the same proper mass $\mu$, while the $i^{\mathrm{th}}$ layer has a MSH mass $\mu_{i}$ for $i=0,\ldots,N$, where $i=0$ refers to the innermost core, and $i=N$ labels the outermost shell. The cumulative mass $M_{i}$ inside the $i^{\rm th}$ layer is then given by
@@
The differences between our description of a ball of dust particles and other approaches to the quantisation of a dust ball are discussed in details in Ref. [1]. We just remark that our procedure amounts to quantising the trajectories of individual dust particles like is done for the electron in the quantum mechanics of the hydrogen atom, whereas canonical approaches like that in Ref. [3] quantise collective degrees of freedom, such as the radius of the ball or of shells. The relativistic action for geodesics of dust particles is their proper time (times the mass $\mu$), which yields the mass-shell condition in the form of the Hamiltonian constraint (1.3) [4]. The canonical quantisation prescription therefore leads to a time-independent Schrödinger equation,
$\hat{H}_{i}\,\psi_{n_{i}}=\mathcal{E}_{i}\,\psi_{n_{i}}\ ,$ (1.4)
whose solutions are given by the Hamiltonian eigenfunctions $\psi_{n_{i}}$ corresponding to the eigenvalues
@@
We also recall that the measure in this Hilbert space is chosen to ensure that $\hat{H}_{i}$ are Hermitian (precisely like for the hydrogen atom) [1]. Since $E_{i}^{2}\geq 0$, Eq. (1.3) sets the bound $n_{i}\geq N_{i}$, where  [^2] See also Ref. [5].
@@
with $L^{1}_{n-1}$ the generalised Laguerre polynomials and $n=1,2,\ldots$. Finally, the wave-functions (1.7) are normalised in the scalar product which makes $\hat{H}_{i}$ Hermitian:
@@
This mass function weakens the central singularity of the Schwarzschild solution and matches the total ADM mass at the surface of the outermost layer, as shown in the left panel of Fig. 1. Eq. (1.12) thus provides an effective description of the geometry of a collapsed core of radius $R_{\rm s}\simeq\bar{R}_{N}+\Delta{R}_{N}\simeq 3\,G_{\rm N}\,M/2$, shorter than, but still comparable to, the gravitational radius $R_{\rm H}=2\,G_{\rm N}\,M$ (more details about the effective energy-momentum tensor of the ground state and interior geometry can be found in Ref. [1]).
@@ 2.1 Quantum mass refinement
The wave-function (1.7) of each layer shows an oscillatory damped pattern that suggests an overlap among (at least) the nearest neighbours. This effect turns the linear mass profile into a (slightly) parabolic mass distribution inside the ball (all the details can be found in Ref. [9]). In fact, particles from the $i^{\rm{th}}$ layer have non-vanishing probability to be located inside the $j^{\rm{th}}$ layer given by Eq. (2.2) (see right panel of Fig. 1).
@@
with $x=r/R_{H}$ and $a$, $b$, $c$ fitting coefficients [9].  [^3] The profile does not significantly depend on the number $N$ of layers since the largest faction of the total mass is located in the outer layers. This refined distribution preserves the integrability of the metric in the origin and does not give rise to any inner horizons. Values of $\mathcal{M}_{i}$ and the parabolic profile are shown in the left panel of Fig. 2.
@@ 3 Quasi-normal modes
where the parameter $\epsilon$ keeps track of the expansion order. Once (3.13) is replaced into the master equation (3.9)  [^4] The second order term in Eq. (3.9) is to be considered multiplied by $\epsilon^{2}$. one obtains the expressions for the $S_{n}$. The first two terms $S_{0}$ and $S_{1}$ provide a solution of the form
@@
and the next goal is to determine the matrix elements $M_{ij}$. The final step consists into matching WKB solutions in region I and III with the one in region II. This is done by Taylor expanding $Q(r_{*})$ around its maximum and solve equation (3.9) with a combination of parabolic cylinder functions. The coefficients of these asymptotically expanded solutions, can be matched with those in the outer region and used to populate the matrix (3.15). The boundary conditions for $\xi$ at $r_{*}\to\pm\infty$ follow from requiring that no waves can escape from the horizon, i.e. $A^{(\mathrm{in})}_{\rm H}=0$, and that there are no incoming waves, i.e. $A^{(\mathrm{in})}_{\infty}=0$.  [^5] These arguments are reversed if $\omega_{R}<0$. These conditions on the amplitudes forces the respective coefficients $M_{ij}$ to vanish, which finally allows us to derive the expression
@@
We also remark that the calculations for the interpolating case are performed to fifth order for practical considerations, since higher orders of Padé approximants become more computationally demanding than advantageous. For some values of $n$ and $\ell$,  [^6] Specifically when $n\geq\ell$ where the method is less accurate. the polynomials involve very small numbers that compromise the numerical stability of their asymptotic convergence, rather than improving the precision. This is not a real drawback, as fifth-order frequencies are very accurate and there is no a priori reason to force higher-order corrections [16].
@@ 4 Conclusion and outlooks
However, the linear mass function does not account for the full quantum nature of the dust particles in the collapsed core. In particular, it discards the overlapping of the eigenfunctions (1.7) shown in the right panel of Fig. 1, whose squared amplitude weighs the location of the particles inside each layer and yields the parabolic mass distribution [9] shown in Eq. (2.5) and the left panel of Fig. 2. In this work, we further refined the description of the outermost layer, to account for the quantum leaking of the dust particles into the region outside the horizon. This is the effect that can lead to modifications of (linear) perturbations in the exterior space. To show that neglecting this quantum leakage reproduces the Schwarzschild phenomenology, we also considered the mass function (2.9) that interpolates smoothly between the interior linear MSH mass and the total ADM mass at a value of the areal coordinate $r=R_{\rm s}<R_{\rm H}$.
@@
Data Availability Statement
Data sharing was not applicable to this article as no data sets were generated or analyzed in this study.
Conflict of Interest
The authors declare no conflict of interest.